Suppose an MNC has $120 million worth of euro accounts receivables and $90 million worth of euro accounts payables. The spot rate is currently at $1.1234/€. How much is the actual exposure of the MNC to the euro.  Please think it through before you answer and only after studying. “I thought you meant this or I thought you meant that” will not be a valid excuse. You should know exactly what I mean and only one answer is correct.
Answer #1

When the multinational corporation will be having similar amount of receivables and payables in foreign currency during the similar period of time we can net the payments with the receipts and we can arrive at our net figure to calculate the exposure.

Total dollars receivables are= 120 Millions

Total dollars payable are= 30 Millions

Total net exposure in $=[receivables in $ - payables in$]

=(120-90)= 30 Millions Dollars.

Converting these dollars to the euros

= (30 *1.1234)

= Euros 33.702